A scale reads 398 N when a piece of copper is hanging from it. What does it read (in N) when it is lowered so that the copper is submerged in water?


1
Expert's answer
2021-04-19T17:05:51-0400

"mass_{copper} = w\/g=398\/9.8=40.6122kg"

"density \space of \space copper=8960kg/m^3"

"V_{copper}=mass/density=40.6122/8960=4.5326\times 10^{-3}" "m^3"

upthrust force "=density\space of \space water \times volume \times gravity=1000\times4.5326\times10^{-3}\times 9.8=44.4195N"


scale reading"=398N-44.4195N=353.58N"